Left on together the hot and cold water fill up the tub in 4minutes. Cold water takes 14 minutes by itself. How many minutes does it the hot water to fill up by itself?
Answer by lwsshak3(11628)

Left on together the hot and cold water fill up the tub in 4minutes. Cold water takes 14 minutes by itself. How many minutes does it the hot water to fill up by itself?
**
1/14=cold water work rate
1/x=hot water work rate
1/4=work rate with both hot and cold water left on
sum of individual work rates=work rate working together
1/14+1/x=1/4
1/x=1/4-1/14
=14/56-4/56
=10/56
=5/28
x=28/5
ans:
minutes it takes the hot water to fill up by itself=28/5=5.6